OpenStudy (anonymous):

Help please! Instructions: tell whether each equation represents a direct variation. If so, identify the constant of variation. 1). Y=4x+1 2). Y=1/3x 3). 6y=12x Thanks!

OpenStudy (anonymous):

1) y=4x+1 _______NO since y depends on more than one minterm 2) 2) y=1/3x _______NO since x is on the bottom so Inverse variation 3) 6y = 12x _______YES y = kx here where k is constant of variation = 12/6 = 2 P.S ~ ** This is 7th grade homework because I had this last week!

OpenStudy (anonymous):

If you need more help on the next homework, message me because I'll tell the answers the one I have the same. **Goodluck!